OpenStudy (hockeychick23):

A set of y-values is transformed (in order to yield a better linear fit) by taking the natural logarithm of each value. The regression of ln y on x is then computed as ln y=-3.1+2.5x. What is the predicted value of y when x=3?

OpenStudy (loser66):

just replace x =3 in ln y = .... then take e both sides to get y

OpenStudy (hockeychick23):

@iGreen i got 4.4 could you check my answer please?

OpenStudy (hockeychick23):

@rational can you help me with this please?

OpenStudy (rational):

\[\ln y = 4.4\] \[y=?\]

OpenStudy (hockeychick23):

i looked up how to do it and by the steps this is what i got: y=e^2.5x-3.1= (e^2.5)^xe^-3.1= EXP(2.5)^xEXP(-3.1) i just don't know what to do next

OpenStudy (hockeychick23):

and I'm not sure if 4.4 is right that was the solution that i thought might be right

OpenStudy (rational):

\[\ln y =4.4 \] is same as \[y=e^{4.4}\]

OpenStudy (rational):

plug that into calculator

OpenStudy (rational):

http://www.wolframalpha.com/input/?i=e%5E4.4

OpenStudy (hockeychick23):

oh so its 81.45? I'm not sure if the 4.4 i got was right to start though

OpenStudy (rational):

Correct. lny = 4.4 is also correct

OpenStudy (hockeychick23):

oh so the answer is 81.45, thanks! (sorry I'm sorta bad with this)

OpenStudy (rational):

remember : \[\ln y = a\] is same as \[y=e^a\] where \(e\) is an irrational number : \(2.718\ldots\)

OpenStudy (hockeychick23):

oh ok does that mean I'm wrong?

OpenStudy (rational):

you're right, why do u think you're wrong

OpenStudy (hockeychick23):

i thought i had to do something with the 2.718 sorry

OpenStudy (rational):

wolfram did that for you

OpenStudy (rational):

wolfram knows that e = 2.718...

OpenStudy (hockeychick23):

ok awesome thanks!